OpenStudy (reganmckay32):

120(ln(root{5}{x^3}) + ln(root{3}{y^2}) - ln (root{4}{16z^5}))

satellite73 (satellite73):

is that 120 out front?

satellite73 (satellite73):

\[120(\ln(\sqrt[5]{x^3}) +\ln(\sqrt[3]{y^2}) - \ln (\sqrt[4]{16z^5})\]

OpenStudy (reganmckay32):

Correct! Sorry. I just got back on.